(38) A large sample of Northern Pike caught at Taltson Lake (Canada) showed that the average length was x = 32.5 inches with sample standard deviation s = 8.6 inches. (a) Compute the z-score of 46, 21.5 and 17 inch Northern Pikes. (b) Are any of the Northern Pikes from part (a) unusual in size?.
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case you're satisfied with the answer.  

xbar = 32.5

s = 8.6

a.

Z score of 45 is (Xbar-Mu)/s = (32.5-45)/8.6 = -1.454

Similarly, Z scores for 21.5 and 17 are:

Z score of 21.5 = (21.5-32.5)/8.6 = -1.28

Z score of 17 = (21.5-17)/8.6 = 0.52

So, if northern pikes have to unusual then their Z scores have to more than the magnitude of 2, which means >2 or <-2

But that' not the case with any of the value' Z score, hence, none of the Northern Pikes from (a) are unusual in size.